OpenStudy (anonymous):

Find the solution of the square root of the quantity of x plus 3 plus 6 equals 9, and determine if it is an extraneous solution MEDAL AND FAN!!!

OpenStudy (anonymous):

ITS PLUS 6 AFTER THE PLUS 3

OpenStudy (mayankdevnani):

your Question :- \[\huge \sf \sqrt{x+3+6}=9\]

OpenStudy (mayankdevnani):

correct ? @lastkingymcmb

OpenStudy (anonymous):

yes and determine if it is an extraneous solution

OpenStudy (mayankdevnani):

just square both sides,you get \[\huge \sf x+9=81\] Understood so far?

OpenStudy (anonymous):

yes I understand so far

OpenStudy (mayankdevnani):

now,find the value of x?

OpenStudy (anonymous):

which is 72 ? correct

OpenStudy (mayankdevnani):

correct

OpenStudy (anonymous):

plug 72 in for x ??

OpenStudy (mayankdevnani):

yeah

OpenStudy (mayankdevnani):

you'll get answer as 9=R.H.S

OpenStudy (anonymous):

whats R.H.S

OpenStudy (mayankdevnani):

right hand side

OpenStudy (anonymous):

I don't have the right calculator